# Analytic Continual Test-Time Adaptation for Multi-Modality Corruption

### v1

@@ 3.4 Soft Pseudo-label Strategy
Due to the definition of TTA, all datasets (i.e., and ) prior to timestamp are not accessible when solving the optimization problem. However, with the aid of the memory bank , the solution can still be computed, as stated in Theorem 2.
Theorem 2. The optimal solution to Formula 12 is given as
where the memory bank is updated in a recursive way with timestamp , as
The proof of Theorem 2 is also provided in Appendix A. It can be seen that, the size of the memory bank is constant and depends on the dimension of the feature expansion layer and class number . The equations above demonstrate that the size of remains unchanged during adaptation, regardless of the timestamp .
4 Experiments
In this section, we evaluate the proposed MDAA approach under the challenging MM-CTTA setting. The MM-CTTA setting is detailed in Sec. 4.1. Sec. 4.2 compares MDAA with SOTA methods through extensive experiments to demonstrate its superior performance. Additionally, the ablation studies on each component of MDAA are presented in Sec. 4.3, which illustrate the effectiveness of MDAA in addressing different challenges within the MM-CTTA setting. The implementation details are provided in Appendix C.
4.1 Benchmarks and Settings
In this section, we introduce the datasets and task settings used for MM-CTTA. The MM-CTTA setting requires the model to initially train on uncorrupted source datasets. Subsequently, the model performs TTA on each corrupted target domain in sequence. We utilize two datasets for this setting: Kinetics50 (Kay et al., 2017) and VGGSound (Chen et al., 2020). While the original uncorrupted datasets serve as source, the corrupted target datasets, Kinetics50-C and VGGSound-C, are constructed following Yang et al. (2024), which introduces 15 types of video corruptions and 6 audio corruptions at severity level 5.
To evaluate the model performance, we designed two classification tasks specifically for MM-CTTA following previous research on corruption-related TTA (Wang et al., 2022; 2021; Yang et al., 2024; Lei & Pernkopf, 2024). The first task, named progressive single-modality corruption, sequentially introduces different types of corruption to one modality while keeping the other modality uncorrupted. Focusing on evaluating the model’s resistance to catastrophic forgetting, this task is set in an online manner, where the model processes only one sample at a time. The second task, called interleaved modality corruption, continually alternates corruption between the two modalities. While most methods perform poorly in the online setting due to severe catastrophic forgetting, this task uses a batch size of 64 during test time to emphasize assessing the model’s ability to adapt to dynamic reliability biases.
4.2 Performance Comparison
To provide a comprehensive comparison, we reproduce different types of TTA methods under the MM-CTTA setting. Typical TTA methods of TENT (Wang et al., 2021) and SAR (Niu et al., 2023); CTTA methods of CoTTA (Wang et al., 2022) and EATA (Niu et al., 2022); and MM-TTA methods of MMTTA (Shin et al., 2022) and READ (Xiong et al., 2024). To ensure a fair comparison, all methods are based on the pre-trained ViT-baesed CAV-MAE (Gong et al., 2022b) as the multi-modal encoders. When reproduce methods that update the BN layers, we instead update the LN layers to suit the ViT structure. Additionally, we evaluate the Source model as a strong baseline, which is only trained on the source dataset and remains frozen during test time.
Performances of different methods on progressive single-modality corruption are listed in Table 1, 2 and 3, where Table 1 is based on the audio corruption, Table 2 and 3are performance on the video corruption in Kinetics50-C and VGGSound-C. The proposed MDAA achieves SOTA performance on most target domains. It is noteworthy that nearly all comparison models collapse and perform significantly worse than non-updating models on VGGSound-C. In contrast, EATA achieves better results by restricting model parameter updates, which facilitates successful adaptation. Specifically, MDAA outperforms previous methods by 3.00%-3.57% and 3.03%-6.22% on average for audio and video tasks in Kinetics50-C, and by 0.94%-1.10% and 0.13%-0.18% on average for audio and video tasks in VGGSound-C. Furthermore, MDAA consistently maintains its lead over other methods in the later stages of the tasks. These results demonstrate that MDAA is more robust against catastrophic forgetting in MM-CTTA.
@@ 4.2 Performance Comparison
The comparison results of interleaved modality corruption tasks are shown in Table 4 and 5. In this task, EATA, which is more good at memorisation, is not dominant in the task of highlighting reliability bias. READ, which is specifically designed to address intra-modal reliability bias, demonstrates strong performance in this area. However, its effectiveness is limited to video corruption in MM-CTTA, as performance drops significantly during audio corruption. In contrast, MDAA is well-adapted to the corruption of different modalities, outperforming READ by 2.39%-6.84% and EATA by 0.60%-7.28% on average across Kinetics50-C and VGGSound-C.
@@
4.3 Ablation Study
@@
In this section, we conduct three ablation studies on both video-corrupted Kinetics50-C dataset and audio-corrupted VGGSound-C dataset in severity level 5, with the batch size of 64. For simplicity, in the following section we use KS-video and VGG-audio to represent these two tasks.
Component analysis. To verify the effectiveness of each MDAA component, we adopt an ablation study w.r.t three components as shown in Table 6. As observed, the model using only AC underperforms, with an average accuracy 0.5%-0.57% lower than READ on KS-video, and the model even collapses on VGG-audio, with an average accuracy of 0.47%-0.56%. This occurs because, while AC can prevent model forgetting, it cannot filter out unreliable samples, leading to issues with error accumulation and reliability bias. In contrast, the addition of DSM significantly improves performance, with gains of 1.73%-7.52% on KS-video and 33.74%-33.90% on VGG-audio. The introduction of SPS allows the model to learn from more possibly correct labels at the same time, thus further improving the performances on most tasks.
@@
| MDAA (ACs) | 61.82 | 62.02 | 0.47 | 0.56 |
| MDAA (ACs+DSM) | 63.55 | 69.54 | 34.87 | 34.85 |
| MDAA (ACs+DSM+SPS) | 65.43 | 69.30 | 35.82 | 35.77 |
Reliable selection threshold. To examine the effect of the threshold on the DSM, we plot the model’s performance with of 0, 1-4, 5-4, 1-3, 2-3 and 5-3 in Fig.4(A-B). The performance of the model on both datasets exhibits an increasing and then decreasing trend. When is close to , the ACs are updated for nearly every sample, introducing more error. Conversely, when increases too much, the ACs do not learn from new inputs since no samples can pass through the DSM, leading to a decline in performance. Therefore, the we choose 1-3 a moderate valuefor both datasets as the threshold.
Soft label reconstruction. In this part we determine a suitable number to reconstruct the pseudo-labels. Given the sorted top distribution set ,we assign weights to each class in a decreasing manner using the formula . The results of the model are plotted in Fig.4(C-D) with chosen from to . While using soft labels can inevitably introduce error, there are instances where performance with one-hot labels may exceed that of the SPS (e.g., KS-video ). However, SPS remains beneficial as the dataset becomes larger and more complex. In fact, performance using SPS surpasses that of one-hot labels in most tasks, making SPS a worthwhile trade-off in MM-CTTA. Generally, the optimal number for reconstructions depends on the number of categories in the dataset, and a larger number of categories warrants a larger . However, more classes will be included in the reconstruction as increase, thereby introducing more error and reducing model performance. Therefore in SPS, we determine to use top and classes to reconstructed label in Kinetics50-C and VGGSound-C respectively.
5 Conclusion
@@
In this paper, we analysed the factors that affect the model in the MM-CTTA task (i.e., error accumulation, catastrophic forgetting and reliability bias) and demonstrate that typical TTA methods are not suitable for the MM-CTTA task. To address the impact of these factors, we propose a new paradigm MDAA that introduce analytic learning to TTA for the first time. Instead of just adapting the model to the target domain, MDAA integrates the target domain into source domain, and thus prevent network from forgetting. With the help of DSM and SPS, model is able to dynamically and comprehensively process the information provided by each modality and use reliable samples to update. In the future, we will try to adapt this paradigm to more modalities to solve more challenging problems in real scenarios.
Appendix A Proofs of Theorems
In this section, we provide comprehensive proofs of Theorems 1 and 2 presented in the main paper.
Note above equation as , derive for as
Therefore .
Proofs of Theorem 2. Similar to the proof of Theorem 1, we note the optimal formula as and derive it in terms of as
Therefore .
Appendix B Pseudo-code for MDAA

### v2

@@ 3.3. Model Inference and Analytic Classifier Adaptation
Due to the definition of TTA, all datasets (i.e., and ) prior to timestamp are not accessible when solving the optimization problem. However, with the aid of the memory bank , the solution can still be computed, as stated follow.
Proof of Theorem 1 is provided in Appendix D. It can be seen that, only features and labels that are accessible in phase are involved in the adaptation, which fits the definition of TTA. Solution in Eq. 8 is the global optimal solution for all the samples that have appeared, therefore it is not affected by batch size and timestamp, thus solving the catastrophic forgetting at the root.
3.4. Dynamic Late Fusion Mechanism
Although AC-based adaptation effectively mitigates the forgetting issue, it is worth noting that employing the least squares loss as the objective function may be less robust to outliers compared to the cross-entropy loss. This vulnerability increases the risk of model collapse when adapting with corrupted modality inputs. This raises a critical question: how can we reliably select trustworthy samples for updating each AC?
To tackle this challenge, we propose a novel fusion strategy termed Dynamic Late Fusion Mechanism (DLFM). For each input sample, DLFM first identifies the AC yielding the highest MAP score, designating it as the Leader. During adaptation, each AC is updated independently based on its MAP score relative to the Leader. All possible relationships and update rules can be enumerated as four scenarios (see Figure 3).
(i). Close Distributions: When the MAP scores of the AC and the Leader are nearly identical and correspond to the same label, the AC remains unchanged to avoid reinforcing existing knowledge and causing class imbalance.
(ii). Different Labels with Close Distributions: If the MAP scores are similar but correspond to different labels, no update is made to avoid introducing potential errors due to uncertainty in the Leader’s prediction.
(iii). Evenly Distributed Probabilities: If both the AC and the Leader show evenly distributed probabilities with no clear preference, the AC is not updated to prevent adjustments based on uncertain predictions.
(iv). Significant Difference: If the Leader’s MAP is significantly higher than the AC’s, the AC is updated to reflect the Leader’s confident prediction.
To summarize, cases (i),(ii) and (iii) all belong to a small gap between MAP scores of AC and Leader, while case (iv) belongs to a larger gap. Therefore, given a pre-defined threshold , Rules of DLFM can be noted as:
DLFM leverages the complementary and consistent information across modalities to enhance adaptation. For complementarity, it selects the most reliable prediction—typically from an uncorrupted modality—as a pseudo label to update other classifiers, aligning corrupted inputs with reliable targets to handle domain shifts. For consistency, when all classifiers predict the same label, indicating high confidence and low likelihood of corruption, no update is performed to avoid unnecessary adaptation and prevent class imbalance. Since our recursive formulation already preserves prior knowledge, additional updates are unnecessary. By updating each classifier individually, DLFM maximizes modality-specific information while reducing reliability bias.
It is also important to note that in Eq. 3.3 we treat source data and target data separately, we adjust the category balance for the source dataset by assigning weights to each class. In contrast, it is challenging to apply this approach to the target dataset, as it is impossible to know the exact number of samples in each category within the target domain. However, with the help of the DLFM, category balance can still be maintained, as unimportant samples have been filtered out. Given that the weights added to the sample data are expected to average out to 1, the impact of each sample from both the source and target domains on the model can be considered equivalent.
@@ 3.4. Dynamic Late Fusion Mechanism
4. Experiments
In this section, we evaluate MDAA under two challenging settings which is detailed in Sec. 4.1. Sec. 4.2 compares MDAA with other SOTA methods through extensive experiments and Sec.4.3 conducts ablation studies on MDAA’s components and computational efficiency. Implementation details are in Appendix B.
4.1. Benchmarks and Settings
To evaluate the model performance, we designed two tasks specifically for MM-CTTA. The first task named Progressive single-modality corruption, sequentially corrupts one modality while keeping the other clean, assessing the model’s resistance to catastrophic forgetting. It follows an online setting, where the model processes one sample at a time, leading to an extreme long learning phase. The second task, called interleaved modality corruption, continually alternates corruption between the two modalities. While most methods perform poorly in the online setting due to severe catastrophic forgetting, this task uses a batch size of 64 during test time to emphasize assessing the model’s ability to adapt to dynamic reliability biases. We utilize two datasets for both tasks: Kinetics50 (Kay et al., 2017) and VGGSound (Chen et al., 2020). Following (Yang et al., 2024), we construct their corrupted type named Kinetics50-C and VGGSound-C, each dataset includes 15 types of video corruptions and 6 audio corruptions at the largest severity (level 5).
4.2. Performance Comparison
In this section, we reproduce various TTA methods under the MM-CTTA setting. These include typical TTA methods (TENT (Wang et al., 2021), SAR (Niu et al., 2023)), CTTA methods (EATA (Niu et al., 2022)), and MM-TTA methods (MMTTA (Shin et al., 2022), READ (Xiong et al., 2024)). Additionally, we introduce an MM-CTTA variant, READ+SRM, which integrates READ with the stochastic restoration mechanism (SRM) from CoTTA (Wang et al., 2022). To ensure fairness, all methods use the same pre-trained model CAV-MAE (Gong et al., 2022b) as the multi-modal encoder. For methods that update BN layers, we instead update LN layers to align with the ViT structure of encoder. We also evaluate the source model, a baseline trained only on the source dataset and kept frozen during testing.
The performances of various methods on progressive single-modality corruption are summarized in Tables 1, 2, and 3. Table 1 reports results on audio corruption, while Tables 2 and 3 focus on video corruption. Notably, most comparison methods collapse on VGGSound-C, performing worse than source models. In contrast, EATA achieves better results by limiting parameter updates, enabling more stable adaptation. MDAA outperforms prior methods by 3.00%-3.57% (audio) and 3.03%-6.22% (video) on Kinetics50-C, and by 0.94%-1.10% (audio) and 0.13%-0.18% (video) on VGGSound-C, on average. While MDAA exhibits suboptimal performance in scenarios where data corruption minimally impacts source model performance—attributable to its reduced parameter space for adaptation compared to baselines—it nevertheless demonstrates superior overall adaptability. MDAA consistently retains its advantage in later stages of the tasks, demonstrating superior robustness against catastrophic forgetting in MM-CTTA.
The comparison results of interleaved modality corruption tasks are shown in Table 4 and 5. In this task, EATA, which is more good at memorization, is not dominant in the task of highlighting reliability bias. READ, which is specifically designed to address intra-modal reliability bias, demonstrates strong performance in this area. However, its effectiveness is limited to video corruption in MM-CTTA, as performance drops significantly during audio corruption. In contrast, MDAA is well-adapted to the corruption of different modalities, outperforming READ by 2.39%-6.84% and EATA by 0.60%-7.28% on average across Kinetics50-C and VGGSound-C.
4.3. Ablation Study
@@
All studies in this section use the video-corrupted Kinetics50-C (KS-video) and audio-corrupted VGGSound-C (VGG-audio) datasets at severity level 5 with a batch size of 64.
Component analysis. To verify the effectiveness of each MDAA component, we adopt an ablation study w.r.t three components as shown in Table 6. As observed, due to the vulnerability of least square loss, the model using only ACs underperforms, with an average accuracy 0.5%-0.57% lower than READ (using cross-entropy loss) on KS-video. The model even collapses on VGG-audio, with an average accuracy of 0.47%-0.56%. The introduction of DLFM consistently improves model performance across all types of corruption, achieving the gains of 1.73%-7.52% on KS-video and 33.74%-33.90% on VGG-audio. Soft labeling introduces a small amount of uncertainty compared to hard labeling, it effectively alleviates the negative impact caused by incorrect pseudo labels during TTA. Although such uncertainty may lead to slightly lower performance in some tasks where corruptions in these tasks caused little effect on the model classification, soft labeling represents a favorable trade-off, particularly when dealing with highly complex datasets such as VGGSound. Furthermore, combining DLFM with soft labels yields the best overall performance.
Reliable selection threshold. To examine the effect of the threshold on the DLFM, we plot the model’s performance with of from 1-5 to 1-1 in Figure 4(A-B). The performance of the model on both datasets exhibits an increasing and then decreasing trend. When is close to , the ACs are updated for nearly every sample, introducing more error. Conversely, when increases too much, the ACs do not learn from new inputs since no samples can pass through the DLFM, leading to a decline in performance. Therefore, value of should not be too small, otherwise the model will not be able to filter out the update samples with low execution. For both datasets in the experiment, a reasonable value of is around 1-3.
@@
| MDAA (ACs+Hard) | 61.82 | 62.02 | 0.47 | 0.56 |
| MDAA (ACs+Soft) | 61.33 | 67.47 | 32.50 | 33.00 |
| MDAA (ACs+DLFM+Hard) | 63.55 | 69.54 | 34.87 | 34.85 |
| MDAA (ACs+DLFM+Soft) | 65.43 | 69.30 | 35.82 | 35.77 |
| Method | KS-video | VGG-audio | ||
| time(s) | memory(MB) | time(s) | memory(MB) | |
| TENT* | 1085 | 37620 | 2571 | 37622 |
| SAR | 2149 | 37620 | 4823 | 39220 |
| EATA | 889 | 67286 | 2313 | 38300 |
| MMTTA* | 2574 | 71732 | 5966 | 68840 |
| READ | 1174 | 9176 | 2611 | 9178 |
| MDAA | 1024 | 8784 | 1682 | 8750 |
Soft label reconstruction. Figure 4(C-D) shows the model results with ranging from to . While using soft labels inevitably introduces error, there are cases such as KS-video where one-hot labels may outperform soft labels. However, using soft pseudo-labels remains beneficial as datasets become larger and more complex. The choice of needs to be moderate, as larger datasets with more categories tend to require larger , while more error is likely to be introduced as increases. In this work, we determine to use top and classes for label reconstruction in Kinetics50-C and VGGSound-C respectively.
Runtime and VRAM usage. To demonstrate MDAA’s practicality, we compare its runtime and VRAM usage against various methods in Table 7. MDAA achieves superior efficiency with lower memory overhead and faster execution, thanks to its unique design. Unlike other methods, MDAA eliminates backpropagation, reducing gradient computation and optimizer storage costs. The primary computational cost of MDAA lies in matrix inversion, with a time complexity of , where is the expanded feature dimension, fixed at 8000 in our experiments.
5. Conclusion
@@
MM-CTTA shows great potential in real-world scenarios, as it enables pre-trained models to handle multi-modal inputs and adapt to continuously changing target domains. However, it also suffers from critical challenges such as reliability bias and catastrophic forgetting. To tackle these issues, we propose a new paradigm, MDAA, which introduces analytic learning into TTA. Instead of simply adapting the model to the target domain, MDAA integrates the target domain into source domain, and thus prevent network from forgetting. With the help of DLFM, model is able to dynamically and comprehensively process the information provided by each modality and use reliable samples to update. Moreover, compared to most SOTA TTA methods, MDAA achieves faster computational times and lower memory usage, highlighting its robust applicability in real-time deployment. In the future, we will try to adapt this method to more modalities to solve more challenging tasks.
Acknowledgements.
This research was supported by the National Natural Science Foundation of China (Grant No. 62306117), the Guangzhou Basic and Applied Basic Research Foundation (2024A04J3681), and GJYC program of Guangzhou (2024D03J0005).References
- (1)
- Boudiaf et al. (2022) Malik Boudiaf, Romain Mueller, Ismail Ben Ayed, and Luca Bertinetto. 2022. Parameter-free online test-time adaptation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8344–8353.
- Cao et al. (2023) Haozhi Cao, Yuecong Xu, Jianfei Yang, Pengyu Yin, Shenghai Yuan, and Lihua Xie. 2023. Multi-modal continual test-time adaptation for 3d semantic segmentation. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 18809–18819.
- Chen et al. (2019) Chaoqi Chen, Weiping Xie, Wenbing Huang, Yu Rong, Xinghao Ding, Yue Huang, Tingyang Xu, and Junzhou Huang. 2019. Progressive feature alignment for unsupervised domain adaptation. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 627–636.
- Chen et al. (2020) Honglie Chen, Weidi Xie, Andrea Vedaldi, and Andrew Zisserman. 2020. Vggsound: A large-scale audio-visual dataset. In ICASSP 2020-2020 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 721–725.
- Dosovitskiy et al. (2020) Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, et al. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. In International Conference on Learning Representations.
- Fang et al. (2024) Di Fang, Yinan Zhu, Runze Fang, Cen Chen, Ziqian Zeng, and Huiping Zhuang. 2024. AIR: Analytic Imbalance Rectifier for Continual Learning. arXiv preprint arXiv:2408.10349 (2024).
- Gao et al. (2023) Jin Gao, Jialing Zhang, Xihui Liu, Trevor Darrell, Evan Shelhamer, and Dequan Wang. 2023. Back to the source: Diffusion-driven adaptation to test-time corruption. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 11786–11796.
- Gong et al. (2022a) Taesik Gong, Jongheon Jeong, Taewon Kim, Yewon Kim, Jinwoo Shin, and Sung-Ju Lee. 2022a. Note: Robust continual test-time adaptation against temporal correlation. Advances in Neural Information Processing Systems 35 (2022), 27253–27266.
- Gong et al. (2022b) Yuan Gong, Andrew Rouditchenko, Alexander H Liu, David Harwath, Leonid Karlinsky, Hilde Kuehne, and James Glass. 2022b. Contrastive audio-visual masked autoencoder. arXiv preprint arXiv:2210.07839 (2022).
- Guo and Lyu (2004) Ping Guo and Michael R Lyu. 2004. A pseudoinverse learning algorithm for feedforward neural networks with stacked generalization applications to software reliability growth data. Neurocomputing 56 (2004), 101–121.
- Guo and Jin (2025) Zirun Guo and Tao Jin. 2025. Smoothing the shift: Towards stable test-time adaptation under complex multimodal noises. arXiv preprint arXiv:2503.02616 (2025).
- Hinton et al. (2015) Geoffrey Hinton, Oriol Vinyals, and Jeff Dean. 2015. Distilling the Knowledge in a Neural Network. stat 1050 (2015), 9.
- Kay et al. (2017) Will Kay, Joao Carreira, Karen Simonyan, Brian Zhang, Chloe Hillier, Sudheendra Vijayanarasimhan, Fabio Viola, Tim Green, Trevor Back, Paul Natsev, et al. 2017. The kinetics human action video dataset. arXiv preprint arXiv:1705.06950 (2017).
- Lei and Pernkopf (2024) Jixiang Lei and Franz Pernkopf. 2024. Two-Level Test-Time Adaptation in Multimodal Learning. In ICML 2024 Workshop on Foundation Models in the Wild.
- Liang et al. (2024) Jian Liang, Ran He, and Tieniu Tan. 2024. A comprehensive survey on test-time adaptation under distribution shifts. International Journal of Computer Vision (2024), 1–34.
- McCloskey and Cohen (1989) Michael McCloskey and Neal J Cohen. 1989. Catastrophic interference in connectionist networks: The sequential learning problem. In Psychology of learning and motivation. Vol. 24. Elsevier, 109–165.
- Müller et al. (2019) Rafael Müller, Simon Kornblith, and Geoffrey E Hinton. 2019. When does label smoothing help? Advances in neural information processing systems 32 (2019).
- Niu et al. (2022) Shuaicheng Niu, Jiaxiang Wu, Yifan Zhang, Yaofo Chen, Shijian Zheng, Peilin Zhao, and Mingkui Tan. 2022. Efficient test-time model adaptation without forgetting. In International conference on machine learning. PMLR, 16888–16905.
- Niu et al. (2023) Shuaicheng Niu, Jiaxiang Wu, Yifan Zhang, Zhiquan Wen, Yaofo Chen, Peilin Zhao, and Mingkui Tan. 2023. Towards Stable Test-time Adaptation in Dynamic Wild World. In The Eleventh International Conference on Learning Representations.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International conference on machine learning. PMLR, 8748–8763.
- Shin et al. (2022) Inkyu Shin, Yi-Hsuan Tsai, Bingbing Zhuang, Samuel Schulter, Buyu Liu, Sparsh Garg, In So Kweon, and Kuk-Jin Yoon. 2022. Mm-tta: multi-modal test-time adaptation for 3d semantic segmentation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 16928–16937.
- Vaswani (2017) A Vaswani. 2017. Attention is all you need. Advances in Neural Information Processing Systems (2017).
- Wang et al. (2021) Dequan Wang, Evan Shelhamer, Shaoteng Liu, Bruno Olshausen, and Trevor Darrell. 2021. Tent: Fully Test-Time Adaptation by Entropy Minimization. In International Conference on Learning Representations.
- Wang et al. (2022) Qin Wang, Olga Fink, Luc Van Gool, and Dengxin Dai. 2022. Continual test-time domain adaptation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 7201–7211.
- Wang et al. (2020) Weiyao Wang, Du Tran, and Matt Feiszli. 2020. What makes training multi-modal classification networks hard?. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 12695–12705.
- Xiong et al. (2024) Baochen Xiong, Xiaoshan Yang, Yaguang Song, Yaowei Wang, and Changsheng Xu. 2024. Modality-Collaborative Test-Time Adaptation for Action Recognition. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 26732–26741.
- Yang et al. (2024) Mouxing Yang, Yunfan Li, Changqing Zhang, Peng Hu, and Xi Peng. 2024. Test-time Adaptation against Multi-modal Reliability Bias. In The Twelfth International Conference on Learning Representations.
- Yang and Hu (2008) Shuang-Hong Yang and Bao-Gang Hu. 2008. A stagewise least square loss function for classification. In Proceedings of the 2008 SIAM International Conference on Data Mining. SIAM, 120–131.
- Zhang et al. (2015) Xu Zhang, Felix Xinnan Yu, Shih-Fu Chang, and Shengjin Wang. 2015. Deep transfer network: Unsupervised domain adaptation. arXiv preprint arXiv:1503.00591 (2015).
- Zhang et al. (2023) Yifan Zhang, Xue Wang, Kexin Jin, Kun Yuan, Zhang Zhang, Liang Wang, Rong Jin, and Tieniu Tan. 2023. Adanpc: Exploring non-parametric classifier for test-time adaptation. In International Conference on Machine Learning. PMLR, 41647–41676.
- Zhao et al. (2025) Yusheng Zhao, Junyu Luo, Xiao Luo, Jinsheng Huang, Jingyang Yuan, Zhiping Xiao, and Ming Zhang. 2025. Attention Bootstrapping for Multi-Modal Test-Time Adaptation. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 39. 22849–22857.
- Zhou et al. (2023) Zhi Zhou, Lan-Zhe Guo, Lin-Han Jia, Dingchu Zhang, and Yu-Feng Li. 2023. ODS: Test-Time Adaptation in the Presence of Open-World Data Shift. In Proceedings of the 40th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 202), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.). PMLR, 42574–42588.
- Zhuang et al. (2021) Huiping Zhuang, Zhiping Lin, and Kar-Ann Toh. 2021. Blockwise recursive Moore–Penrose inverse for network learning. IEEE Transactions on Systems, Man, and Cybernetics: Systems 52, 5 (2021), 3237–3250.
- Zhuang et al. (2023) Huiping Zhuang, Zhenyu Weng, Run He, Zhiping Lin, and Ziqian Zeng. 2023. Gkeal: Gaussian kernel embedded analytic learning for few-shot class incremental task. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 7746–7755.
- Zhuang et al. (2022) Huiping Zhuang, Zhenyu Weng, Hongxin Wei, Renchunzi Xie, Kar-Ann Toh, and Zhiping Lin. 2022. ACIL: Analytic class-incremental learning with absolute memorization and privacy protection. Advances in Neural Information Processing Systems 35 (2022), 11602–11614.
Appendix A Pseudo-code for MDAA